Determine the slope and y-intercept of the two graphs: 3x+y =8 and -x+2y=4
Answer by lwsshak3(11628)

Determine the slope and y-intercept of the two graphs: 3x+y =8 and -x+2y=4
..
Given equations are straight lines. Standard form for straight lines: y=mx+b, m=slope and b=y-intercept. Convert given equations to this form to fine slope and y-intercept
..
3x+y=8
y=-3x+8
slope=-3
y-intercept=8
..
-x+2y=4
2y=x+4
y=x/2 + 4
slope=1/2
y-intercept=4